Do NOT remove anything from this box except in accordance with the Boxer Rebellion rites and the authorization of the Skinner! (At least bubble in any books you take and notify Panthercomm.)

Page 1  The MITSFS Pinkdex  BP39
ERR  TITLE  AUTHOR
≈ M  Artifact Of Evil [@Greyhawk �2] ∘  Gygax, Gary
≈ M  Night Arrant [Gord The Rogue] ∘  Gygax, Gary
≈ M  Saga Of Old City [@Greyhawk �1] ∘  Gygax, Gary
≈ M  Sea Of Death [Gord The Rogue] ∘  Gygax, Gary
≈ M  Joy Wagon, The ∘  Hadley, Arthur T.
≈ M  Planet Big Zero ∘  Hadley, Franklin
≈ M  King Solomon’s Mines [Allan Quatermain] ∘  Haggard, H. Rider
≈ M  People Of The Mist, The ∘  Haggard, H. Rider
≈ M  She [Ayesha 1] ∘  Haggard, H. Rider
≈ M  Mutants Are Coming, The ∘  Haiblum, Isidore
≈ M  Transfer To Yesterday ∘  Haiblum, Isidore
≈ M  Tsaddik Of The Seven Wonders, The ∘  Haiblum, Isidore
≈ M  Wilk Are Among Us, The ∘  Haiblum, Isidore
≈ M  Perry’s Planet [Star Trek] ∘  Haldeman, Jack C. Ii
≈ M  Vector Analysis ∘  Haldeman, Jack C. Ii
≈ M  All My Sins Remembered ∘  Haldeman, Joe
≈ M  Buying Time ∘  Haldeman, Joe
≈ M  Camou�age ∘  Haldeman, Joe
≈ M  Coming, The ∘  Haldeman, Joe
≈ M  Dealing In Futures ∘  Haldeman, Joe
≈ M  Forever Free ∘  Haldeman, Joe
≈ M  Forever Peace ∘  Haldeman, Joe
≈ M  Forever War, The ∘  Haldeman, Joe
≈ M  Hemingway Hoax, The ∘  Haldeman, Joe
≈ M  In�nite Dreams ∘  Haldeman, Joe
≈ M  Mindbridge ∘  Haldeman, Joe
≈ M  Planet Of Judgment [Star Trek] ∘  Haldeman, Joe
≈ M  Tool Of The Trade ∘  Haldeman, Joe
≈ M  World Without End [Star Trek] ∘  Haldeman, Joe
≈ M  Worlds Apart [Worlds 2] ∘  Haldeman, Joe
≈ M  Worlds Enough And Time [Worlds 3] ∘  Haldeman, Joe
≈ M  Blind Spot, The ∘  Hall, Austin/Flint, Homer Eon
≈ M  Armies Of Daylight, The [Darwath 3] ∘  Hambly, Barbara
≈ M  Bride Of The Rat God ∘  Hambly, Barbara
≈ M  Children Of The Jedi [Star Wars] ∘  Hambly, Barbara
≈ M  Crossroad [Star Trek] ∘  Hambly, Barbara
≈ M  Dark Hand Of Magic, The [Sunwolf And Starhawk 3] ∘  Hambly, Barbara
≈ M  Dog Wizard [Windrose 3] ∘  Hambly, Barbara
≈ M  Dragonsbane [Winterlands 1] ∘  Hambly, Barbara
≈ M  Ghost-Walker [Star Trek] ∘  Hambly, Barbara
≈ M  Icefalcon’s Quest [Darwath 5] ∘  Hambly, Barbara
≈ M  Ishmael [Star Trek] ∘  Hambly, Barbara
≈ M  Ladies Of Mandrigyn, The [Sunwolf And Starhawk 1] ∘  Hambly, Barbara
≈ M  Mother Of Winter [Darwath 4] ∘  Hambly, Barbara
≈ M  Rainbow Abyss, The [Sun-Cross 1] ∘  Hambly, Barbara
≈ M  Silent Tower, The [Windrose 1] ∘  Hambly, Barbara
≈ M  Silicon Mage, The [Windrose 2] ∘  Hambly, Barbara
≈ M  Song Of Orpheus [Beauty And The Beast] ∘  Hambly, Barbara
≈ M  Stranger At The Wedding ∘  Hambly, Barbara
≈ M  Those Who Hunt The Night [London Vampires 1] ∘  Hambly, Barbara

Page 2  The MITSFS Pinkdex  BP39
ERR  TITLE  AUTHOR
≈ M  Time Of The Dark, The [Darwath 1] ∘  Hambly, Barbara
≈ M  Traveling With The Dead [London Vampires 2] ∘  Hambly, Barbara
≈ M  Walls Of Air, The [Darwath 2] ∘  Hambly, Barbara
≈ M  Witches Of Wenshar, The [Sunwolf And Starhawk 2] ∘  Hambly, Barbara
≈ M  Battle For The Stars ∘  Hamilton, Edmond
≈ M  Best Of Edmond Hamilton, The ∘  Hamilton, Edmond
≈ M  City At World’s End ∘  Hamilton, Edmond
≈ M  Crashing Suns ∘  Hamilton, Edmond
≈ M  Doomstar ∘  Hamilton, Edmond